### v1

# Order Is All You Need for Categorical Data Clustering

Abstract.
Categorical data composed of nominal valued attributes are ubiquitous in knowledge discovery and data mining tasks. Due to the lack of well-defined metric space, categorical data distributions are difficult to intuitively understand. Clustering is a popular technique suitable for data analysis. However, the success of clustering often relies on reasonable distance metrics, which happens to be what categorical data naturally lack. Therefore, the cluster analysis of categorical data is considered a critical but challenging problem. This paper introduces the new finding that the order relation among attribute values is the decisive factor in clustering accuracy, and is also the key to understanding the categorical data clusters. To automatically obtain the orders, we propose a new learning paradigm that allows joint learning of clusters and the orders. It turns out that clustering with order learning achieves superior clustering accuracy, and the learned orders provide intuition for understanding the cluster distribution of categorical data. Extensive experiments with statistical evidence and case studies have verified the effectiveness of the new “order is all you need” insight and the proposed method. The source code of this submission is temporarily opened [^1] https://anonymous.4open.science/r/OCL-demo.
Cluster analysis, Categorical data, Order relationship, Distance learning, Partitional clustering
submissionid: 51ccs: Computing methodologies Cluster analysisccs: Computing methodologies Learning latent representationsccs: Computing methodologies Mixture modeling
1. Introduction
[…]
In the literature, there are two main types of clustering algorithms for categorical data: 1) $k$-modes and its subsequent variants (Huang, 1998)(Bai et al., 2011)(Jia and Cheung, 2018) that partition data samples into clusters based on certain distance measures, and 2) cluster quality assessment-based algorithms (Hu et al., 2022)(Li et al., 2004) that adjust the sample-cluster affiliations based on certain cluster quality measures. Their performance is commonly dominated by the metric space represented by the adopted measures. Thus, the focus of subsequent research has shifted from designing clustering algorithms to proposing reasonable measures that suit the categorical data clustering (Naouali et al., 2020).
[…]
The conventional Hamming distance measure and one-hot encoding strategy are very common in categorical data clustering, but they both simply represent the distance between two attribute values based on whether the two values are identical or not, which is far from informative. Later, on the one hand, more advanced distance measures have been proposed to discriminatively indicate the value-level differences by considering the intra-attribute probability of possible values (Sandler and Lindenbaum, 2011), inter-attribute relationship (Ahmad and Dey, 2007)(Le and Ho, 2005)(Ienco et al., 2012), or both (Jia et al., 2016)(Zhang and Cheung, 2022b). On the other hand, advanced representation strategies that encode distances among data samples (Qian et al., 2015), and couplings in the attribute and value levels (Jian et al., 2018c)(Jian et al., 2017), have also been proposed Although all the above advanced techniques have considerably improved the clustering performance, they are all task-independent and thus incompetent in adapting different clustering tasks on various categorical datasets.
[…]
It is noteworthy that order is the actual basis for clustering in most cases. For instance, when a $k$-means-type algorithm tries to partition a sample into one of the $k$ clusters, it actually obtains the overall order of the $k$ clusters in terms of admitting the sample. The overall order is formed based on the similarity orders of the clusters to the sample reflected by different attributes. Note that each numerical attribute corresponds to a one-dimensional Euclidean distance space with a clear order of values, which makes it easy to obtain the value order. According to the above analysis, we also note that although most existing categorical data clustering approaches do not specify the order of attribute values, their clustering process still passively forms the order through similarity measurement, regardless of whether there is an obvious semantic order of the attribute values.
Therefore, knowing the correct order of attribute values may be beneficial to categorical data clustering, and the experimental results in Figure 1 confirmed this. It can be seen that even randomly generated orders can easily achieve significant accuracy improvements. Another interesting observation is that, although attribute values of some datasets (i.e., the right eight datasets) are with obvious semantic orders, they are rarely optimal for clustering. This is because the semantic orders of attributes are relatively independent of specific clustering tasks, and this information gap can prevent some semantic orders from well-serving clustering. Therefore, obtaining the optimal orders w.r.t. certain clustering tasks is the key to achieving accurate categorical data clustering.
This paper, therefore, proposes a new clustering paradigm that dynamically learns orders based on the current data partition during clustering. Since the order learning is connected to the clustering task, the optimal order for clustering can thus be automatically obtained. To make full use of the learned orders, a distance learning mechanism is also introduced to dynamically measure the sample-cluster affiliation under the guidance of the orders. It turns out that the learned orders can significantly improve clustering accuracy and can also intuitively help understand the distribution and clustering effect of categorical data. Extensive experimental evaluations demonstrate the superiority of the proposed method. The main contributions can be summarized into three points:
  • A new insight that “obtaining the optimal orders is the decisive factor for accurately clustering categorical data” is proposed for the first time. Accordingly, a new clustering paradigm with order distance learning is proposed.
  • We propose a distance structure learning mechanism to automatically obtain the order and distance. This learning process is jointly performed with clustering, making the clustering performance more robust to various datasets.
  • Both the learning process and the learned order are highly interpretable, which is helpful for understanding complex categorical data. Moreover, the algorithm is efficient, strictly convergent, and can be easily extended to mixed data [^2] Dataset composed of a mixture of categorical and numerical attributes..
2. Related Work
This section overviews the related works in data encoding, representation learning, and consideration of value order, for categorical data clustering.
The data encoding techniques convert the categorical data into numerical values based on certain strategies, and then perform clustering on the numerical encoding. The conventional one-hot encoding and Hamming distance construct similar distance structures for categorical data by assigning distances “1” and “0” to any pair of different and identical values, respectively. As they cannot distinguish different dissimilarity degrees (dos Santos and Zárate, 2015)(Boriah et al., 2008) of data samples, information loss is usually unavoidable during clustering. Some statistic-based distance measures (Lin, 1998)(Cheung and Jia, 2013)(dos Santos and Zárate, 2015) consider the intra-attribute occurrence frequency of values to better distinguish the data samples. They take a common principle that two similar values may have more consistent statistics within the corresponding attribute or cluster. Some metrics (Ahmad and Dey, 2007)(Le and Ho, 2005) further define distance considering the relevance between attributes. The representative measure proposed in (Ienco et al., 2012) selects more relevant attributes for the distance measurement. Later, the measures (Jia et al., 2016)(Jian et al., 2018a) that more comprehensively consider both the intra- and inter-attribute couplings reflected by the data statistics have been presented. Coupling-based encoding strategies (Jian et al., 2018c)(Jian et al., 2017) have also been designed accordingly to encode the couplings. All the above techniques perform value-level encoding to reflect sample-level distances, while the work (Qian et al., 2015) directly encodes a data sample by concatenating the distances from it to all the other samples.
Recently, more advanced representation learning methods have been proposed to learn the categorical data encoding. The work proposed in (Zhu et al., 2022) projects dataset through different kernels for more informative coupling representation, and then learns the kernel parameters to obtain task-oriented representation. Some approaches (Jian et al., 2018b)(Zhu et al., 2020) have also been designed to simultaneously encode the numerical and categorical attributes. However, the involved hyper-parameters have a great impact on their performance, and the appropriate parameter selection will be laborious in different real applications. Although a parameter-free learning approach (Jia and Cheung, 2018) has been proposed to learn sample-cluster distance and number of clusters simultaneously, the distance is updated dependent on cluster statistics and thus cannot reveal the natural distance structure of categorical data.
Some new advances proposed with the consideration of value order further distinguish categorical attributes into nominal and ordinal attributes. The distance metrics (Zhang et al., 2020)(Zhang and Cheung, 2022b) define the distance between attribute values based on the data statistical information with considering the semantic order. Later, a representation learning approach was introduced by (Zhang and Cheung, 2020), which treats the possible values of an ordinal attribute as linearly arranged concepts, and learns the distances between adjacent concepts. Later, the works (Zhang and Cheung, 2022a)(Zhang and Cheung, 2023) further unify the distances of nominal and ordinal attributes, and make them learnable with clustering. The most recent work (Zhang et al., 2022) attempts to convert the nominal attributes into ordinal ones through geometric projection, and then learn the latent distance spaces during clustering. The above recent advances achieve considerable clustering performance improvements and empirically prove the importance of value order in categorical data clustering. However, they all assume the availability of semantic order of attribute values. Moreover, they directly adopt this order for clustering, and overlook that the semantic order may not be suitable under different clustering scenarios as shown in Figure 1.
3. Proposed Method
This section first formulates the problem, then proposes the order distance definition and order learning mechanism, and finally introduces the joint learning of the orders and clusters. Analyses of time complexity and convergence are also provided.
3.1. Problem Formulation
The problem of categorical data clustering with order learning is formulated as follows. Given a categorical dataset $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{n}\}$ with $n$ data samples (e.g., a collection of $n$ bank clients). Each data sample $\mathbf{x}_{i}$ can be denoted as a $d$-dimensional row vector $\mathbf{x}_{i}=[x_{i,1},x_{i,2},..,x_{i,d}]^{\top}$ represented by $d$ attributes $A=\{\mathbf{a}_{1},\mathbf{a}_{2},...,\mathbf{a}_{d}\}$ (e.g., occupation, gender, priority, etc.). Each attribute $\mathbf{a}_{r}$ (e.g., occupation) can be denoted as a column vector $\mathbf{a}_{r}=[x_{1,r},x_{2,r},...,x_{n,r}]$ composed of the occupation description of all the $n$ clients. For the categorical attribute “occupation” $\mathbf{a}_{r}$, all its $n$ values are taken from a limited number of qualitative possible values $V_{r}=\{v_{r,1},v_{r,2},...,v_{r,l_{r}}\}$, e.g., $\{\text{lawyer},\text{doctor},...,\text{scientist}\}$. The subscripts indicate the sequential number of an attribute and the sequential number of the attribute’s possible value. For example, doctor is the second value in the $r$-th attribute “occupation”, so we have $v_{r,2}=\text{doctor}$. We use $l_{r}$ to indicate the total number of possible values of an attribute, and it usually satisfies $l_{r}\ll n$ for real categorical data. Please note that the sequential numbers of possible values only distinguish them, and do not indicate their order.
Assume that there is an optimal order indicating an appropriate underlying distance structure of possible values w.r.t. a clustering task. We denote the optimal order of an attribute $\mathbf{a}_{r}$ as $O^{*}_{r}=\{o^{*}(v_{r,1}),o^{*}(v_{r,2}),...,o^{*}(v_{r,l_{r}})\}$ where the superscript “*” marks the optimum and the value of $o^{*}(v_{r,g})$ is an integer reflecting the unique ranking of $v_{r,g}$ among all the $l_{r}$ possible values in $V_{r}$. By still taking the “occupation” attribute as an example, if its three possible values $v_{r,1}=\text{lawyer}$, $v_{r,2}=\text{doctor}$, and $v_{r,3}=\text{scientist}$, ranking second (i.e., $o(\text{lawyer})=2$), third (i.e., $o(\text{doctor})=3$), and first (i.e., $o(\text{scientist})=1$), respectively. Then the corresponding order of these values are $O_{r}=\{o(\text{lawyer}),o(\text{doctor}),o(\text{scientist})\}=\{2,3,1\}$. Such order reflects a distance structure that roughly satisfies $\text{dist}(v_{r,g},v_{r,h})\propto|o(v_{r,g})-o(v_{r,h})|$. More specifically, the following distance relationship should hold:
  1. (1)
    $\text{dist}(\text{doctor},\text{scientist})>\text{dist}(\text{doctor},\text{lawyer})$,
  2. (2)
    $\text{dist}(\text{doctor},\text{scientist})>\text{dist}(\text{lawyer},\text{scientist})$,
because “doctor” and “scientist” are further away in order, compared to “lawyer” and “doctor” that are adjacent in order.
We aim to approximate the optimal orders of all the $d$ attributes $O^{*}=\{O^{*}_{1},O^{*}_{2},...,O^{*}_{d}\}$. Then the order information reflecting the distance structure can guide the partition of data objects during clustering. Our approach, i.e., learning order relations, is more generative than most existing approaches that directly define the distance between attribute values. We focus more on how the intrinsic relationship between values affects the distance between samples in forming clusters, rather than simply defining the clustering-irrelevant distance between values based on data statistics, e.g., occurrence probability of values in an attribute.
For a crisp partitional clustering task to partition $X$ into $k$ non-overlapping subsets $C=\{C_{1},C_{2},...,C_{k}\}$, the objective can be formalized as the problem of minimizing:
(1) $L(\mathbf{Q},O)=\sum_{m=1}^{k}\sum_{i=1}^{n}q_{i,m}\cdot\Theta(\mathbf{x}_{i},C_{m};O)$
where $\mathbf{Q}$ is an $n\times k$ matrix with its $(i,m)$-th entry $q_{i,m}$ indicating the affiliation between sample $\mathbf{x}_{i}$ and cluster $C_{m}$. Specifically, $q_{i,m}=1$ indicates that $\mathbf{x}_{i}$ belongs to $C_{m}$, while $q_{i,m}=0$ indicates that $\mathbf{x}_{i}$ belongs to a cluster other than $C_{m}$, which can be written as:
(2) $q_{i,m}=\left\{\begin{array}[]{ll}1,&\text{if}\ m=\arg\min\limits_{y}\Theta(\mathbf{x}_{i},C_{y};O)\\ 0,&\text{otherwise}\\ \end{array}\right.$
with $\sum_{i=m}^{k}q_{i,m}=1$. The distance $\Theta(\mathbf{x}_{i},C_{m};O)$ reflects the dissimilarity between $\mathbf{x}_{i}$ and $C_{m}$ computed based on the distance structures reflected by the orders $O$. Therefore, the problem to be solved can be specified as how to jointly learn $\mathbf{Q}$ and $O$ to minimize $L$. When the algorithm converges, a sub-optimal solution with $k$ compact bank client clusters is expected to be obtained. In general, “compact” means that the bank clients with more similar characteristics (i.e., with more similar value descriptions across all the attributes) are gathered into the same cluster, and different clusters are relatively distinguishable.
Frequently used symbols in this paper and the corresponding explanations are sorted out in Table 1 below. Then we discuss how to define the order distance and learn order for accurate categorical data clustering.
Table 1. Frequently used symbols. Note that we uniformly use lowercase, uppercase, bold lowercase, and bold uppercase to indicate value, set, vector, and matrix, respectively.
Symbol Explanation
$X$, $A$, $O$, and $C$ Dataset, attribute set, order set, and cluster set
$\mathbf{x}_{i}$ and $x_{i,r}$ $i$-th data object and $r$-th value of $\mathbf{x}_{i}$
$\mathbf{a}_{r}$, $V_{r}$, and $v_{r,g}$ $r$-th attribute, its value set, and $g$-th value of $V_{r}$
$l_{r}$ Total number of possible values in $V_{r}$
$O_{r}$ Order values (integers) corresponding to $V_{r}$
$o(v_{r,g})$ Order value (integer) of possible value $v_{r,g}$
$\mathbf{Q}$ and $C_{m}$ Object-cluster affiliation matrix and $m$-th cluster
$q_{i,m}$ $(i,m)$-th entry of $\mathbf{Q}$ indicating affiliation of $\mathbf{x}_{i}$ to $C_{m}$
$\Theta(\mathbf{x}_{i},C_{m};O)$ Overall order distance between $\mathbf{x}_{i}$ and $C_{m}$
$\theta(x_{i,r},\textbf{p}_{m,r})$ Order distance between $\mathbf{x}_{i}$ and $C_{m}$ reflected by $\mathbf{a}_{r}$
$\mathbf{D}$ and $\mathbf{d}_{i,r}$ Order distance matrix and its $(i,r)$-th entry
$d_{i,r,g}$ Order distance between $x_{i,r}$ and $v_{r,g}$ ($g$-th value of $\mathbf{d}_{i,r}$)
$\mathbf{P}$ and $\mathbf{p}_{m,r}$ Conditional probability matrix and its $(m,r)$-th entry
$p_{m,r,g}$ Occurrence probability of $v_{r,g}$ in $C_{m}$ ($g$-th value of $\mathbf{p}_{m,r}$)
$O_{r,m}$ Order values of $O_{r}$ obtained through $C_{m}$
$L_{r,m}$ A fraction of $L$ jointly contributed by $\mathbf{a}_{r}$ and $C_{m}$
3.2. Order Distance
We first define the sample-cluster distance $\Theta(\mathbf{x}_{i},C_{m};O)$ with considering the order $O$, to connect the order with the clustering objective. The distance $\Theta(\mathbf{x}_{i},C_{m};O)$ is defined as:
(3) $\Theta(\mathbf{x}_{i},C_{m};O)=\frac{1}{d}\sum_{r=1}^{d}\theta(x_{i,r},\textbf{p}_{m,r})$
where
(4) $\theta(x_{i,r},\textbf{p}_{m,r})=\mathbf{d}_{i,r}^{\top}\mathbf{p}_{m,r}.$
is the order distance between $\mathbf{x}_{i}$ and all the samples in $C_{m}$ reflected by the $r$-th attribute. $\mathbf{d}_{i,r}$ is an $l_{r}$-dimensional vector containing the order distances between $x_{i,r}$ and possible values in $V_{r}$, and $\mathbf{p}_{m,r}$ is an $l_{r}$-dimensional vector describing the probability distribution of $V_{r}$ within cluster $C_{m}$, which can be written as:
(5) $\mathbf{d}_{i,r}=[d_{i,r,1},d_{i,r,2},...,d_{i,r,l_{r}}]$
and
(6) $\mathbf{p}_{m,r}=[p_{m,r,1},p_{m,r,2},...,p_{m,r,l_{r}}],$
respectively. The $g$-th value of $\mathbf{d}_{i,r}$ is defined as the normalized order distance between $x_{i,r}$ and $v_{r,g}$, which can be written as:
(7) $d_{i,r,g}=\frac{|o(x_{i,r})-o(v_{r,g})|}{l_{r}-1}$
where $l_{r}-1$ in the denominator is to normalize the order distance into the interval $[0,1]$ to ensure distance comparability of different attributes. The $g$-th value of $\mathbf{p}_{m,r}$ is the occurrence probability of $v_{r,g}$ in cluster $C_{m}$, which can be written as:
(8) $p_{m,r,g}=\frac{card(X_{r,g}\cap C_{m})}{card(C_{m})}$
where $X_{r,g}=\{\mathbf{x}_{i}|x_{i,r}=v_{r,g}\}$ is a sample set collecting all the samples in $X$ with their $r$-th values equal to $v_{r,g}$, and the cardinality function $card(\cdot)$ counts the number of elements in a set.
Remark 1.
Rationality of the order distance. The intra-cluster probability distribution and the value order are introduced to form the order distance. The probability descriptor $\mathbf{p}_{m,r}$ is adopted to informatively represent the value distribution in different clusters for appropriate distance measurement and accurate order learning in Section 3.3. By making the distance $\mathbf{d}_{i,r}$ dependent on the order $O$ through Eq. (7), we facilitate a direct connection between the order learning and the clustering tasks. Such a connection leverages the optimization of the order $O$ for more accurate clustering in Section 3.4.
3.3. Order Learning
Based on the above-defined order distance, we discuss how to obtain the optimal order:
(9) $O^{*}=\arg\min_{O}L(\mathbf{Q},O).$
A simple but extremely laborious solution is to search different order combinations across all the attributes. There are a total of $\prod_{r=1}^{d}l_{r}!$ combinations, and each combination should be evaluated on $n$ samples by accumulating their sample-cluster distances calculated by Eq. (3), which makes the search infeasible. A relaxed efficient solution is to search the optimal order within clusters by assuming independence of attributes. Accordingly, optimal order of $\mathbf{a}_{r}$ w.r.t. $C_{m}$ denoted as $O^{*}_{r,m}=\{o^{*}_{m}(v_{r,1}),o^{*}_{m}(v_{r,2}),...,o^{*}_{m}(v_{r,l_{r}})\}$ can be obtained through the efficient intra-cluster traversal search:
(10) $O^{*}_{r,m}=\arg\min_{O_{r}}L_{r,m}$
where $L_{r,m}$ is an inner sum component of the objective $L$. More specifically, $L_{r,m}$ is jointly contributed by the $r$-th attribute and $m$-th cluster, which can be derived from Eqs. (1) and (3) as:
(11) $L_{r,m}=\sum_{\mathbf{x}_{i}\in C_{m}}\theta(x_{i,r},\textbf{p}_{m,r}).$
By searching for the optimal ranking in each cluster and then performing a linear combination based on the cluster size, the overall ranking of each possible value $v_{r,g}$ can be obtained by:
(12) $\bar{o}(v_{r,g})=\sum_{m=1}^{k}\left(o^{*}_{m}(v_{r,g})\cdot\frac{card(C_{m})}{n}\right).$
Since the value of $\bar{o}(v_{r,g})$ is not necessarily an integer, we sort the obtained rankings $\{\bar{o}(v_{r,1}),\bar{o}(v_{r,2}),...,\bar{o}(v_{r,l_{r}})\}$ of an attribute in ascending order to obtain the final ranking of values in $V_{r}$, and the corresponding integer order $\tilde{O}_{r}=\{\tilde{o}(v_{r,1}),\tilde{o}(v_{r,2}),...,\tilde{o}(v_{r,l_{r}})\}$ is obtained. Correspondingly, we collectively refer to all orders obtained through Eqs. (10) - (12) as $\tilde{O}=\{\tilde{O}_{1},\tilde{O}_{2},...,\tilde{O}_{d}\}$.
Remark 2.
Principle of the order learning. In the process of obtaining $\tilde{O}$, the optimal order of each attribute is determined for each cluster by Eqs. (10) - (12). Through Eq. (11), the specific distribution of samples within each cluster is fully considered (see Remark 1) to appropriately reflect the component $L_{r,m}$ of the objective $L$. Hence, an expected minimization on a series of $L_{r,m}$s can be achieved by the determined orders. The overall order of an attribute is obtained by integrating the contributions of all the clusters through Eq. (12). Here, the cluster sizes are taken into account to accommodate more to larger clusters, thereby achieving a more effective reduction on $L$.
3.4. Joint Order and Cluster Learning
We facilitate joint learning of clusters and orders by integrating the order learning into the clustering objective, which can be rewritten in a more detailed form based on Eqs. (1), (3), and (4):
(13) $L(\mathbf{Q},O)=\sum_{m=1}^{k}\sum_{i=1}^{n}q_{i,m}\cdot\frac{1}{d}\sum_{r=1}^{d}\mathbf{d}_{i,r}^{\top}\mathbf{p}_{m,r}$
where the order distance $\mathbf{D}=\{\mathbf{d}_{i,r}|i\in\{1,2,...,n\},r\in\{1,2,...,d\}\}$ is dependent to the order $O$ according to Eqs. (5) and (7), and the probability distribution $\mathbf{P}=\{\mathbf{p}_{m,r}|m=\{1,2...,k\},r\in\{1,2,...,d\}\}$ is dependent to the partition $\mathbf{Q}$ according to Eqs. (6) and (8). Typically, $\mathbf{Q}$ and $O$ are iteratively computed to minimize $L$ following a $k$-modes-type optimization process. Specifically, given $\hat{O}$, the order distance $\mathbf{D}$ can be updated accordingly, and then a new $\mathbf{Q}$ can be computed by:
(14) $q_{i,m}=\left\{\begin{array}[]{ll}1,&\text{if}\ m=\arg\min\limits_{y}\frac{1}{d}\sum_{r=1}^{d}\mathbf{d}_{i,r}^{\top}\mathbf{p}_{y,r}\\ 0,&\text{otherwise}\\ \end{array}\right.$
with $i=\{1,2,...,n\}$ and $m=\{1,2,...,k\}$. Eq. (14) is a detailed version of Eq. (2) that adopts the order distance defined by Eqs. (3) and (4). Then, $\mathbf{P}$ is updated according to the new $\mathbf{Q}$. When the iterative updating of $\mathbf{Q}$ and $\mathbf{P}$ converges, we obtain $O$ according to Eqs. (10) - (12), and update $\mathbf{D}$ accordingly. In summary, $L$ is minimized by iteratively solving the following two sub-problems:
  1. (1)
    Fix $\hat{\mathbf{Q}}$ and $\hat{\mathbf{P}}$, solve the minimization problem $L(\hat{\mathbf{Q}},O)$ by searching $O$ and updating $\mathbf{D}$ accordingly;
  2. (2)
    Fix $\hat{O}$ and $\hat{\mathbf{D}}$, solve the minimization problem $L(\mathbf{Q},\hat{O})$ by iteratively computing $\mathbf{Q}$, and updating $\mathbf{P}$ according to $\mathbf{Q}$, until convergence;
The whole algorithm named Order and Cluster Learning (OCL) is summarized as Algorithm 1. The next subsection provides analyses of time complexity and convergence to further explain OCL.
Algorithm 1 OCL: Order and Cluster Learning Algorithm
0:  Dataset $X$, number of sought clusters $k$
0:  Partition $\mathbf{Q}$, order $O$
1:  Initialization: Set outer loop counter and convergence mark by $E\leftarrow 0$ and $Conv\_O\leftarrow False$, respectively; Randomly partition $X$ into $k$ clusters $C$, obtain $\mathbf{Q}^{\{I\}}$ and $\mathbf{P}^{\{I\}}$ accordingly; Randomly initialize $O^{\{E\}}$, update $\mathbf{D}^{\{E\}}$ accordingly;
2:  while $Conv\_O=False$ do
3:     $E\leftarrow E+1$; Compute $O^{\{E\}}$ by Eqs. (10) - (12); Update $\mathbf{D}^{\{E\}}$ accordingly; Set inner loop counter and convergence mark by $I\leftarrow 0$ and $Conv\_I\leftarrow False$, respectively;
4:     while $Conv\_I=False$ do
5:        $I\leftarrow I+1$; Compute $\mathbf{Q}^{\{I\}}$ by Eq. (14); Update $\mathbf{P}^{\{I\}}$ accordingly; Obtain $L(\mathbf{Q}^{\{I\}},O^{\{E\}})$;
6:        if $L(\mathbf{Q}^{\{I\}},O^{\{E\}})\geq L(\mathbf{Q}^{\{I-1\}},O^{\{E\}})$ then
7:           $Conv\_I\leftarrow True$; $\mathbf{Q}^{\{E\}}=\mathbf{Q}^{\{I\}}$;
8:        end if
9:     end while
10:     if $L(\mathbf{Q}^{\{E\}},O^{\{E\}})\geq L(\mathbf{Q}^{\{E-1\}},O^{\{E-1\}})$ then
11:        $Conv\_O\leftarrow True$;
12:     end if
13:  end while
3.5. Complexity and Convergence Analysis
We analyze the time complexity to show the scalability of OCL, and then provide its convergence analysis.
Theorem 1.
The time complexity of OCL is $O(EInkd\varsigma+Edkn\varsigma!)$.
Proof. Assume it involves $I$ inner iterations (lines 4 - 9 in Algorithm 1) to compute $\mathbf{Q}$ and $\mathbf{P}$, and $E$ outer iterations (lines 2 - 13 in Algorithm 1) to search $O$ and update $\mathbf{D}$. For worst-case analysis, we adopt $\varsigma$ to indicate the maximum number of possible values of a dataset by $\varsigma=\max(l_{1},l_{2},...,l_{d})$.
For each of the $I$ inner iterations, we obtain $\mathbf{P}$ upon the $n$ data samples with complexity $O(nkd\varsigma)$, and obtain $\mathbf{D}$ upon the $\varsigma$ values of $d$ attributes on $n$ samples with complexity $O(nd\varsigma)$. Then the $n$ samples are partitioned to $k$ clusters with complexity $O(nkd\varsigma)$. For $I$ iterations in total, the complexity is $O(Inkd\varsigma)$.
For each of the $E$ outer iterations, a new order of $\varsigma$ possible values of each attribute within each cluster is searched by Eqs. (10) - (12), and the complexity for $d$ attributes and $k$ clusters is thus $O(dkn\varsigma!)$. For $E$ outer iterations with considering the $I$ inner iterations, the overall time complexity of OCL is $O(EInkd\varsigma+Edkn\varsigma!)$. □
Remark 3.
Scalability of OCL. The essential characteristic of categorical data is that the number of possible values for an attribute is finite and small, usually ranging from 2 to 8 for real benchmark datasets as shown in the Appendix. Accordingly, $\varsigma$ can be treated as a small constant, and the time complexity is simplified to $O(EInkd)$, which is linear to the scaling of $n$ or $d$.
Theorem 2.
OCL algorithm converges to a local minimum in a finite number of iterations.
Proof. We prove the convergence of the inner and outer loop, i.e., lines 4 - 9 and lines 2 - 13 of Algorithm 1, respectively.
Lemma 0.
The inner loop converges to a local minimum of $L(\mathbf{Q},\hat{O})$ in a finite number of iterations with fixed $\hat{O}$. Note that the number of possible partitions $\mathbf{Q}$ is finite for a dataset with a finite number of objects $n$. We then illustrate that each possible partition $\mathbf{Q}$ occurs at most once by OCL. Assume that two optimizers satisfy $\mathbf{Q}^{\{I_{1}\}}=\mathbf{Q}^{\{I_{2}\}}$, where $I_{1}\neq I_{2}$. Then it is clear that $L(\mathbf{Q}^{\{I_{1}\}},O^{\{E\}})=L(\mathbf{Q}^{\{I_{2}\}},O^{\{E\}})$. However, the value sequence of the objective function $\{L(\mathbf{Q}^{\{1\}},O^{\{E\}}),L(\mathbf{Q}^{\{2\}},O^{\{E\}}),...,L(\mathbf{Q}^{\{I\}},O^{\{E\}})\}$ generated by OCL is strictly decreasing. Hence, the result follows.
Lemma 0.
The outer loop converges to a local minimum of $L(\mathbf{Q},O)$ in a finite number of iterations. We note that the status of orders $O$ is in a discrete space and thus the number of possible orders is finite. The number of possible partitions $\mathbf{Q}$ is also finite. Therefore, the number of possible combinations of $\mathbf{Q}$ and $O$ is finite. We then prove that each possible combination of $\mathbf{Q}$ and $O$ appears at most once by OCL. We note that given $\mathbf{Q}^{\{E_{1}\}}=\mathbf{Q}^{\{E_{2}\}}$, where $E_{1}\neq E_{2}$, the corresponding $O^{\{E_{1}+1\}}$ and $O^{\{E_{2}+1\}}$ can be computed accordingly, satisfying $O^{\{E_{1}+1\}}=O^{\{E_{2}+1\}}$. Then we can obtain their corresponding partitions $\mathbf{Q}^{\{E_{1}+1\}}$ and $\mathbf{Q}^{\{E_{2}+1\}}$ at the convergence of the next inner loop, satisfying $\mathbf{Q}^{\{E_{1}+1\}}=\mathbf{Q}^{\{E_{2}+1\}}$. It is clear that $L(\mathbf{Q}^{\{E_{1}+1\}},O^{\{E_{1}+1\}})=L(\mathbf{Q}^{\{E_{2}+1\}},O^{\{E_{2}+1\}})$. However, the sequence $\{L(\mathbf{Q}^{\{1\}},O^{\{1\}}),L(\mathbf{Q}^{\{2\}},O^{\{2\}}),...,L(\mathbf{Q}^{\{E\}},O^{\{E\}})\}$ is monotonically decreasing. Hence, the result follows.
Since both the inner loop and outer loop of Algorithm 1 converge to a local minimum, convergence of OCL can be confirmed. □
4. Experiments
Different experiments have been designed to evaluate the performance of the proposed order learning strategy as well as the corresponding clustering algorithm. Detailed experimental settings and the obtained results are introduced as follows.
Table 2. Statistics of the 12 datasets. $d^{\text{\{nom\}}}$, $d^{\text{\{ord\}}}$, and $n$ are the numbers of nominal attributes, ordinal attributes, and samples, respectively. $k^{*}$ is the number of true clusters.
No. Dataset Abbrev. $d^{\text{\{nom\}}}$ $d^{\text{\{ord\}}}$ $n$ $k^{*}$
1 Soybean (small) SB 35 0 47 4
2 Nursery School NS 1 7 12960 4
3 Amphibians AP 6 6 189 2
4 Inflammations Diagnosis DS 0 5 120 2
5 Caesarian Section CS 1 3 80 2
6 Hayes-Roth HR 2 2 132 3
7 Zoo ZO 15 0 101 7
8 Breast Cancer BC 5 4 286 2
9 Lymphography LG 15 3 148 4
10 Tic-Tac-Toe TT 9 0 958 2
11 Australia Credit AC 0 8 690 2
12 Congressional Voting VT 16 0 435 2
4.1. Experimental Settings
Four types of experiments have been conducted to evaluate the proposed method. A brief outline of the experiments is given below: (1) Clustering performance evaluation with significance test (Section 4.2): We compared the proposed OCL with conventional and state-of-the-art methods on different types of datasets. To statistically analyze the performance differences between variant approaches, we also conducted a significance test on the clustering results; (2) Ablation studies (Section 4.3): To more specifically demonstrate the effectiveness of the core components of OCL and investigate the role of learned order in improving clustering performance, we conduct ablation studies from the algorithm aspect and data aspect. (3) Convergence and efficiency evaluation (Section 4.4): The values of the objective function after each learning iteration have been recorded to study the convergence of OCL. Its efficiency has also been studied with experiments on large datasets with scaling sizes of samples $n$ and attributes $d$. (4) Visualization of clustering effect (Section 4.5): To visually demonstrate the effectiveness of the order distance metric learned by OCL, we use it and the counterparts to numerically encode the data, and utilize the t-SNE for visualization.
Nine counterparts have been selected for the comparative study. The counterparts include traditional K-MoDes (KMD) algorithm (Huang, 1998), popular entropy-based metric LSM (Lin, 1998), and context-based distance measurement CBDM (Ienco et al., 2009). The latter two are representative measures proposed for categorical data, both of which can be embedded into KMD to conduct clustering. Besides, some state-of-the-art methods have also been taken into account, including Jia’s Distance Metric (JDM) (Jia et al., 2016), Unified Distance Metric based Clustering (UDMC) (Zhang and Cheung, 2022b), HD based Clustering (HDC) (Zhang and Cheung, 2022a), projection based clustering Het2Hom (H2H) (Zhang et al., 2022), Distance Learning based Clustering (DLC) algorithm (Zhang and Cheung, 2020), and graph metric based clustering algorithm ADC (Zhang and Cheung, 2023). Please note that we have set the parameters (if any) of these methods to the values recommended in the corresponding paper. In the experiments, each clustering result was recorded based on the average performance of 10 runs.
Three validity indices, i.e., clustering accuracy (CA) (He et al., 2005), ARI coefficients (Rand, 1971)(Gates and Ahn, 2017), and standardized mutual information NMI (Estévez et al., 2009), have been selected to evaluate the clustering performance. Usually, larger values of these indices indicate better clustering performance. We also conduct the Bonferroni Dunn (BD) test [^3] The Friedman test (Friedman, 1937) is performed first to ensure a meaningful BD post-hoc test. (Demšar, 2006) on the performance of different methods and calculate the critical difference (CD) interval to statistically analyze the superiority of our OCL.
[…]
Table 3. Clustering performance evaluated by CA. “$\overline{\textit{AR}}$” row reports the average performance rankings.
Datasets KMD LSM JDM CBDM UDMC DLC H2H HDC ADC OCL
SB 0.8468$\pm$0.15 0.8106$\pm$0.16 0.7894$\pm$0.14 0.7957$\pm$0.15 0.7915$\pm$0.18 0.8000$\pm$0.14 0.9234$\pm$0.12 0.7915$\pm$0.15 0.7660$\pm$0.13 0.9830$\pm$0.05
NS 0.3208$\pm$0.05 0.2975$\pm$0.04 0.2980$\pm$0.04 - 0.3074$\pm$0.05 0.3386$\pm$0.07 0.3113$\pm$0.05 0.3208$\pm$0.05 0.3208$\pm$0.05 0.3573$\pm$0.05
AP 0.5392$\pm$0.03 0.5656$\pm$0.04 0.6159$\pm$0.04 0.5926$\pm$0.03 0.5698$\pm$0.03 0.6032$\pm$0.00 0.5603$\pm$0.03 0.5608$\pm$0.04 0.5529$\pm$0.04 0.6222$\pm$0.00
DS 0.6525$\pm$0.08 0.6525$\pm$0.08 0.7208$\pm$0.12 0.7208$\pm$0.12 0.6867$\pm$0.12 0.6650$\pm$0.12 0.6175$\pm$0.04 0.6867$\pm$0.12 0.6950$\pm$0.12 0.7458$\pm$0.08
CS 0.5462$\pm$0.03 0.5488$\pm$0.06 0.5788$\pm$0.07 0.5875$\pm$0.04 0.5825$\pm$0.05 0.6350$\pm$0.03 0.6075$\pm$0.02 0.5837$\pm$0.05 0.5525$\pm$0.04 0.6313$\pm$0.04
HR 0.3758$\pm$0.02 0.3886$\pm$0.03 0.3917$\pm$0.02 0.3924$\pm$0.04 0.3773$\pm$0.02 0.3697$\pm$0.04 0.3333$\pm$0.00 0.4000$\pm$0.03 0.3758$\pm$0.02 0.4326$\pm$0.02
ZO 0.6624$\pm$0.12 0.6733$\pm$0.11 0.6743$\pm$0.11 0.6752$\pm$0.10 0.6693$\pm$0.10 0.6802$\pm$0.08 0.7059$\pm$0.05 0.6733$\pm$0.10 0.6743$\pm$0.11 0.7792$\pm$0.10
BC 0.5472$\pm$0.04 0.5594$\pm$0.07 0.6392$\pm$0.09 0.6346$\pm$0.09 0.5413$\pm$0.06 0.6140$\pm$0.09 0.5266$\pm$0.06 0.6052$\pm$0.10 0.6185$\pm$0.10 0.6650$\pm$0.05
LG 0.4277$\pm$0.06 0.4446$\pm$0.04 0.4547$\pm$0.06 0.4905$\pm$0.06 0.4466$\pm$0.05 0.5095$\pm$0.08 0.4959$\pm$0.05 0.4959$\pm$0.05 0.5101$\pm$0.04 0.5426$\pm$0.02
TT 0.5569$\pm$0.05 0.5593$\pm$0.04 0.5563$\pm$0.04 0.5408$\pm$0.03 0.5386$\pm$0.03 0.5637$\pm$0.03 0.5640$\pm$0.04 0.5465$\pm$0.04 0.5471$\pm$0.04 0.5785$\pm$0.03
AC 0.7281$\pm$0.12 0.7783$\pm$0.08 0.6335$\pm$0.12 0.6354$\pm$0.12 0.7683$\pm$0.08 0.7488$\pm$0.14 0.7703$\pm$0.08 0.6762$\pm$0.14 0.6293$\pm$0.14 0.8206$\pm$0.08
VT 0.8625$\pm$0.01 0.8699$\pm$0.00 0.8699$\pm$0.00 0.8761$\pm$0.00 0.8667$\pm$0.00 0.8531$\pm$0.09 0.8736$\pm$0.00 0.8736$\pm$0.00 0.8713$\pm$0.00 0.8943$\pm$0.00
$\overline{\textit{AR}}$ 7.58 6.46 5.71 5.13 7.08 4.67 5.33 5.63 6.33 1.08
Table 4. Clustering performance evaluated by ARI (with range [-1,1]). “$\overline{\textit{AR}}$” row reports the average performance rankings.
Datasets KMD LSM JDM CBDM UDMC DLC H2H HDC ADC OCL
SB 0.7638$\pm$0.18 0.7555$\pm$0.19 0.7232$\pm$0.19 0.7407$\pm$0.19 0.7307$\pm$0.22 0.7148$\pm$0.20 0.8650$\pm$0.22 0.7350$\pm$0.19 0.6968$\pm$0.17 0.9620$\pm$0.12
NS 0.0376$\pm$0.03 0.0341$\pm$0.02 0.0283$\pm$0.02 - 0.0393$\pm$0.03 0.0779$\pm$0.08 0.0309$\pm$0.03 0.0376$\pm$0.03 0.0376$\pm$0.03 0.1015$\pm$0.06
AP 0.0022$\pm$0.01 0.0171$\pm$0.02 0.0530$\pm$0.03 0.0322$\pm$0.02 0.0178$\pm$0.02 0.0355$\pm$0.00 0.0076$\pm$0.02 0.0150$\pm$0.02 0.0066$\pm$0.02 0.0524$\pm$0.00
DS 0.1077$\pm$0.13 0.1077$\pm$0.13 0.2347$\pm$0.20 0.2347$\pm$0.20 0.1844$\pm$0.20 0.1556$\pm$0.19 0.0512$\pm$0.05 0.1844$\pm$0.20 0.1921$\pm$0.19 0.2590$\pm$0.16
CS -0.0023$\pm$0.01 0.0088$\pm$0.04 0.0295$\pm$0.04 0.0250$\pm$0.03 0.0248$\pm$0.04 0.0647$\pm$0.03 0.0342$\pm$0.02 0.0234$\pm$0.03 0.0040$\pm$0.02 0.0627$\pm$0.03
HR -0.0081$\pm$0.01 -0.0008$\pm$0.01 -0.0013$\pm$0.01 -0.0010$\pm$0.01 -0.0046$\pm$0.01 -0.0033$\pm$0.01 -0.0149$\pm$0.00 0.0038$\pm$0.02 -0.0064$\pm$0.01 0.0368$\pm$0.01
ZO 0.5787$\pm$0.17 0.6085$\pm$0.17 0.5996$\pm$0.17 0.6091$\pm$0.15 0.6063$\pm$0.16 0.6094$\pm$0.12 0.6442$\pm$0.07 0.6138$\pm$0.15 0.6143$\pm$0.15 0.7536$\pm$0.11
BC 0.0046$\pm$0.02 0.0240$\pm$0.05 0.0811$\pm$0.07 0.0777$\pm$0.07 0.0128$\pm$0.04 0.0523$\pm$0.05 -0.0033$\pm$0.00 0.0605$\pm$0.08 0.0692$\pm$0.08 0.0799$\pm$0.03
LG 0.0843$\pm$0.05 0.0951$\pm$0.03 0.1027$\pm$0.06 0.1432$\pm$0.07 0.1018$\pm$0.04 0.1544$\pm$0.09 0.1715$\pm$0.04 0.1095$\pm$0.06 0.1262$\pm$0.05 0.1552$\pm$0.03
TT 0.0170$\pm$0.03 0.0184$\pm$0.02 0.0181$\pm$0.02 0.0095$\pm$0.01 0.0091$\pm$0.01 0.0156$\pm$0.01 0.0198$\pm$0.02 0.0120$\pm$0.02 0.0124$\pm$0.02 0.0226$\pm$0.02
AC 0.2596$\pm$0.18 0.3336$\pm$0.12 0.1210$\pm$0.15 0.1199$\pm$0.13 0.3092$\pm$0.11 0.3184$\pm$0.22 0.3107$\pm$0.11 0.1913$\pm$0.18 0.1349$\pm$0.18 0.4313$\pm$0.16
VT 0.5247$\pm$0.02 0.5463$\pm$0.01 0.5463$\pm$0.01 0.5648$\pm$0.01 0.5368$\pm$0.01 0.5200$\pm$0.18 0.5572$\pm$0.00 0.5572$\pm$0.00 0.5503$\pm$0.00 0.6207$\pm$0.00
$\overline{\textit{AR}}$ 7.88 5.92 5.50 5.38 6.71 5.08 5.29 5.50 6.42 1.33
4.2. Clustering Performance Evaluation
In this section, we investigate the clustering performance of different algorithms and statistically analyze the superiority of OCL.
4.2.1. Clustering performance on different datasets
The clustering results obtained by different algorithms on various datasets in terms of different validity indices are shown in Tables 3 and 4. The best and second-best results on each dataset are highlighted in bold and underlined, respectively. The observations include the following four aspects: (1) Overall, OCL reaches the best result on most datasets, indicating its superiority in clustering. (2) On the CA index, although OCL does not perform significantly better than the second-best method on the AP and TT datasets, the second-best methods are different on these datasets, but OCL has stable performance on them. (3) On the CA index, although OCL does not perform the best on CS dataset, its performance is close to the best result (the difference is less than 0.005), which still proves OCL’s competitiveness. (4) In terms of the ARI index, there are four datasets on which OCL does not perform the best, namely AP, CS, BC, and LG. However, similarly, the differences between OCL and the best-performing algorithms on AP, CS, and BC datasets are very small. In addition, the result of CBDM is not reported on the NS dataset because all the NS’s attributes are independent of each other, making the inter-attribute dependence-based CBDM fail in measuring distances.
4.2.2. Significance study
We first implement the Friedman (Friedman, 1937) test on the average ranks reported in the last rows in Tables 3 and 4. The corresponding p-values are 0.000017 and 0.000091 respectively, both passing the test under the confidence of 99% (p-value $=0.01$). On this basis, results of the two-tailed BD test (Demšar, 2006) with confidence intervals 0.95 ($\alpha=0.05$) and 0.9 ($\alpha=0.1$) are shown in Figure 2, with the corresponding CD intervals CD_95%=3.4275 and CD_90%=3.1383, respectively. According to the criterion presented in (Demšar, 2006), OCL is considered to perform significantly better than all the counterparts that fall outside the right boundary of the CD intervals. It can be seen that the proposed OCL is significantly superior to all counterparts.
Figure 2. Results of the two-tailed BD tests w.r.t. the CA (a) and ARI (b) performance in Tables 3 and  4, respectively.
4.3. Ablation Studies
To demonstrate the effectiveness of the core components of OCL and to verify our hypothesis on the orders of categorical data attributes, we design several variants of OCL for ablation studies from both algorithm and data perspectives.
From the algorithm aspect, to evaluate the effectiveness of the proposed Order Distance (OD), we compare OCL with its variant OCLI, which uses normalized equidistant order distance without considering the probability distribution of possible values. To evaluate the effectiveness of the Order Learning (OL) mechanism, we further remove the process of order learning from OCLI and only update the order once in the first iteration, which forms another variant OCLII. To completely discard Order Information (OI), we further make OCLII use only the traditional Hamming distance, thus forming the third variant OCLIII. Table 5 compares the clustering performance of OCL and the three variants.
It can be observed that the performance of OCL is superior to all variants. More specifically, OCL outperforms OCLI significantly on 11 out of all 12 datasets. This indicates that OD can effectively utilize the distribution information of the attribute values within different clusters for reasonable order distance measurement. Furthermore, the performance of OCLI is also significantly better than OCLII on all 12 datasets, which verifies the effectiveness of the OL mechanism in searching optimal orders during clustering. In addition, compared to the OCLIII variant that completely disregards OI, OCLII has better performance on 11 out of all 12 datasets, which proves that even a roughly estimated initial order can still significantly and extensively improve clustering performance.
Table 5. CA performance of ablated OCL variants formed by successively ablating the three core components, i.e., Order Distance (OD), Order Learning (OL), and Order Information (OI). The green arrows indicate the performance improvements of an OCL variant over its previous more basic variant. “$\overline{\textit{AR}}$” column reports the average performance rankings.
OCL Components Variants Datasets $\overline{\textit{AR}}$
OD OL OI SB NS AP DS CS HR ZO BC LG TT AC VT
$\checkmark$ $\checkmark$ $\checkmark$ OCL   0.9830$\uparrow$   0.3573$\uparrow$   0.6222$\uparrow$   0.7458$\uparrow$   0.6313   0.4326$\uparrow$   0.7792$\uparrow$   0.6650$\uparrow$   0.5426$\uparrow$   0.5785$\uparrow$   0.8206$\uparrow$   0.8943$\uparrow$   1.17$\uparrow$
$\checkmark$ $\checkmark$ OCLI   0.9149$\uparrow$   0.3364$\uparrow$   0.6085$\uparrow$   0.7283$\uparrow$   0.6500$\uparrow$   0.3765$\uparrow$   0.7525$\uparrow$   0.5643$\uparrow$   0.4872$\uparrow$   0.5784$\uparrow$   0.7546$\uparrow$   0.8942$\uparrow$   2.00$\uparrow$
$\checkmark$ OCLII   0.8681$\uparrow$   0.3359$\uparrow$   0.6069$\uparrow$ 0.  7125$\uparrow$   0.6388$\uparrow$ 0.3712   0.7000$\uparrow$   0.5636$\uparrow$   0.4453$\uparrow$   0.5762$\uparrow$   0.7545$\uparrow$   0.8828$\uparrow$   2.92$\uparrow$
OCLIII 0.8468 0.3208 0.5392 0.6525 0.5463 0.3758 0.6624 0.5472 0.4277 0.5569 0.7281 0.8625 3.92
From the data aspect, to validate our hypothesis that correct order information is crucial to categorical data clustering, we conduct ablation experiments on datasets containing both nominal and ordinal attributes. We compare the proposed OCL, which learns orders for both the nominal and ordinal attributes, to two different variants. The first one is LNRO (Learn Nominal Reserve Ordinal), which retains the semantic order information of the original ordinal attribute and only learns orders for the nominal attributes. The other one is RNRO (Reserve Nominal Reserve Ordinal), which discards the order learning process for both nominal and ordinal attributes. Table 6 compares the clustering performance of the three variants. It can be observed that on five out of the six datasets, the performance of LNRO is superior to RNRO. This verifies our hypothesis that the nominal attributes also contain potential order information that is effective for clustering. Significantly, we can find that the OCL method outperforms LNRO on five out of the six datasets. This indicates that the orders learned by the proposed mechanism for ordinal attributes are more conducive to clustering than the original semantic orders.
Table 6. CA performance of OCL variants treating Nominal Attributes (NA) and Ordinal Attributes (OA) in different ways.
Learn Order for Variants Datasets $\overline{\textit{AR}}$
OA NA NS AP CS HR BC LG
OCL   0.3573$\uparrow$   0.6222$\uparrow$ 0.6313   0.4326$\uparrow$   0.6650$\uparrow$   0.5426$\uparrow$   1.17$\uparrow$
LNRO 0.3281   0.6074$\uparrow$   0.6350$\uparrow$   0.3985$\uparrow$   0.5476$\uparrow$   0.4818$\uparrow$   2.00$\uparrow$
RNRO 0.3308 0.5984 0.6225 0.3902 0.5276 0.4493 2.83
4.4. Convergence and Efficiency Evaluation
To evaluate the convergence of OCL, we execute it on all 12 datasets and present the convergence curves in Figure 3. The horizontal and vertical axes represent the number of learning iterations and the value of the objective function $L$, respectively. In each sub-figure of Figure 3, the blue triangles mark the iterations of order updates, while the red square indicates the iteration that OCL converges. It can be observed that, after each order update, the value of $L$ further decreases. This indicates that the proposed order learning mechanism is consistent with the optimization of the clustering objective function. Moreover, for all the tested datasets, OCL completes the learning process within 30 iterations with at most three updates of the orders. This reflects that OCL converges efficiently with guarantee, which conforms to the analysis in Section 3.5.
Figure 3. Convergence curve of OCL on all 12 datasets.
Furthermore, to evaluate the execution efficiency of OCL, we construct large synthetic datasets with different numbers of samples and attributes to investigate the execution time of OCL in comparison with other counterparts. The results reported in Figure 4 show that the execution time of OCL is better than UDMC and DLC on large-scale datasets, and better than H2H, UDMC, and CBDM on high-dimensional datasets. Additionally, as the number of samples or attributes increases, the execution time of OCL increases linearly, which is consistent with the time complexity analysis in Section 3.5. In summary, OCL is efficient compared to the state-of-the-art methods and does not incur too much additional computational cost compared to the simplest methods.
Figure 4. Execution time on synthetic datasets with different numbers of samples $n$ and attributes $d$.
4.5. Visualization of Clustering Effect
To demonstrate the cluster discrimination ability of the OCL method, we utilize the distance between attribute values learned by OCL, CBDM, and ADC respectively to encode the attributes of the VT dataset. Subsequently, the t-SNE (Maaten and Hinton, 2008) is introduced to process the encoded data into two-dimensional space and visualize it by marking the samples with real labels in different colors. It can be seen from Figure 5 that OCL has significantly better cluster discrimination ability than CBDM and ADC.
(a) CBDM @@@@ VT
(b) ADC @@@@ VT
(c) OCL @@@@ VT
Figure 5. t-SNE visualization of the VT dataset represented by CBDM, ADC, and OCL.
5. Concluding Remarks
This paper introduces a new finding that knowing the correct order relation of attribute values is critical to cluster analysis of categorical data. Accordingly, we propose the OCL paradigm that allows joint learning of data partitions and order relation of attribute values for accurate clustering. The proposed learning mechanism can infer a more appropriate order among categorical attribute values based on the current clusters, thus yielding more compact clusters based on the updated order. The data partition and order learning are iteratively executed to approach the clustering objective, i.e., maximizing the overall similarity of intra-cluster samples. As a result, more accurate clustering results and insightful orders can be obtained simultaneously. OCL converges quickly with guarantee, and extensive experiments conducted on 12 real benchmark datasets fully demonstrate its superiority in comparison with state-of-the-art clustering methods. A case study has also been provided to show the rationality of the learned order in understanding unlabeled categorical data.
The proposed OCL method is also not exempt from limitations. Although it demonstrates the potential to enhance the performance of mixed data clustering, the deep connection between the treatment of categorical and numerical attributes remains to be explored. In addition, it would also be promising to extend OCL to more complex datasets with non-static distribution, an unknown number of clusters, and extremely imbalanced clusters.